In young double slit experiment the nth red bright band coincides with (n+1)th blue bright band. If the wavelength of red and blue lights are 7500A and 5000A  , the value of ‘n is

Grade:12

1 Answers

Arun
6 years ago
Dear Umar
 

Path difference for (n+1)th blue band = (n+1)b

path difference for nth red bright band = nr

(n+1)b= nr

(n+1) 5000= n*7500

2n +2 = 3n

n = 2

 

Regards

Arun (askIITians forum expert)